At a spring concert, tickets for adults cost $4.00 and tickets for students cost $2.50. How many of each kind of tickets were purchase if 125 tickets were bought for $413.00?
Answer by kapilsinghi(68)

Let the number of stubent tickets = x
so number of adult tickets = 125-x
cost of student ticket = 2.5
cost of student tickets = (2.5)x
cost of adult ticket = 4
cost of adult ticket = 4(125-x)
total cost of tickets = (2.5)x + 4(125-x) = 413
(2.5)x + 4(125-x) = 413
(2.5)x + 500 - 4x = 413
500 - 413 = 4x - (2.5)x
87 = (1.5)x
87/1.5 = x
x = 58
number of student tickets = 58
number of adult tickets = 125 - 58 = 67
kapilsinghi123@gmail.com